OpenStudy (anonymous):

If d/dx (f(2x^4))=7x^4 calculate f1(x)

OpenStudy (freckles):

what is the mean of f1?

OpenStudy (freckles):

meaning*

OpenStudy (anonymous):

the derivative

OpenStudy (freckles):

do you know chain rule?

OpenStudy (freckles):

\[\frac{d}{dx}f(2x^4)=(2x^4)' f(2x^4)=7x^4\]

OpenStudy (freckles):

find the derivative of 2x^4